# Adversarial Coevolutionary Illumination with Generational Adversarial MAP-Elites

### v1

@@ Abstract
Unlike traditional optimization algorithms focusing on finding a single optimal solution, Quality-Diversity (QD) algorithms illuminate a search space by finding high-performing solutions that cover a specified behavior space. However, tackling adversarial problems is more challenging due to the behavioral interdependence between opposing sides. Most applications of QD algorithms to these problems evolve only one side, thus reducing illumination coverage. In this paper, we propose a new QD algorithm, Generational Adversarial MAP-Elites (GAME), which coevolves solutions by alternating sides through a sequence of generations. Combining GAME with vision embedding models enables the algorithm to directly work from videos of behaviors instead of handcrafted descriptors. Some key findings are that (1) emerging evolutionary dynamics sometimes resemble an arms race, (2) starting each generation from scratch increases open-endedness, and (3) keeping neutral mutations preserves stepping stones that seem necessary to reach the highest performance. In conclusion, the results demonstrate that GAME can successfully illuminate an adversarial multi-agent game, opening up interesting future directions in understanding the emergence of open-ended coevolution.
@@ Introduction
[…]
Quality Diversity (QD) (Pugh et al.,, 2016) is a branch of evolutionary algorithms that seeks to illuminate a problem, i.e., discovering a diverse set of high-quality solutions that maximize a fitness function while covering a behavior space. It has demonstrated success across various domains, including robotics (Cully et al.,, 2015), procedural content generation (Gravina et al.,, 2019), chemical synthesis (Jiang et al.,, 2022), and aeronautics (Brevault and Balesdent,, 2024).
@@
Several such efforts have already been undertaken. For instance, in the turn-based strategy card game, Hearthstone, Fontaine et al., (2019) illuminate the deck space, while Fontaine et al., (2020) evolve a set of neural network controllers with a fixed deck to find different strategies. More recently, Samvelyan et al., 2024a investigate strategic weaknesses in a Deep Reinforcement Learning (DRL) agent during a simulated football game, while Samvelyan et al., 2024b search for different safety-violating prompts for a Large Language Model (LLM). A limitation of these approaches is that they only optimize one side of the adversarial problem. This is problematic, as adversarial systems often lead to an arms race (Dawkins and Krebs,, 1979), which requires improvements on both sides to emerge.
@@
This paper introduces a new QD algorithm for adversarial problems, which we call Generational Adversarial MAP-Elites (GAME). Combined with a Vision Embedding Model (VEM) such as CLIP (Radford et al.,, 2021), GAME can work from videos of behaviors instead of requiring handcrafted behavior descriptors. GAME handles the embedding space’s high dimensionality by using growing unstructured archives (Vassiliades et al.,, 2017). To evaluate the approach, we apply it to a multi-agent adversarial game and show through ablation studies that it effectively illuminates the space of strategies. The paper’s main contributions are:
  • a new adversarial coevolution QD algorithm, GAME;
  • the use of a VEM as the behavior space in a MAP-Elites algorithm with a growing archive;
  • an empirical evaluation of GAME on a multi-agent adversarial game, including comparisons with ablations.
Related-work
@@
GAME aims to create an open-ended adversarial coevolution of solutions that broaden illumination. Creating an open-ended artificial environment that continually presents new, interesting challenges, and in turn, fosters the evolution of novel and meaningful solutions to those challenges is one of the main quests in the artificial life community (Bedau et al.,, 2000; Dorin and Stepney,, 2024).
@@
A form of non-strictly adversarial problems is the coevolution of an environment and the agent interacting with it. The objective is for the environment to evolve in a way that presents new challenges that are neither too easy nor too difficult, effectively creating a curriculum for the agent by offering a sequence of meaningful stepping stones. Brant and Stanley, (2017) coevolve mazes and maze-solving agents using minimal criterion coevolution, i.e., any agent that solves a maze and any maze that is solved at least once can reproduce. Their approach demonstrates that no handcrafted fitness function or behavior descriptor is required.
@@
In contrast to those works, GAME is not designed for the particular problem of evolving controllers and environments, but applies to any adversarial problem.
@@
Self-play is a closely related field, which creates a learning curriculum by making the agent play against itself or an earlier version of itself. Alpha-Star (Vinyals et al.,, 2019) achieves grandmaster level in the Real-Time Strategy (RTS) player-versus-player game StarCraft 2. It uses a league of agents that serve as training opponents for the main agents. Using a league prevents strategy collapse, improves robustness, and helps avoid local minima and forgetting.
@@
QD algorithms have been used to illuminate adversarial problems. For example, Fontaine et al., (2019) propose MAP-Elites with sliding boundaries to build different decks in Hearthstone, using as opponents first a fixed set of decks and then the best decks found in the first phase to explore the possible counters. Using the six best found decks as opponents, Fontaine et al., (2020) use Covariance Matrix Adaptation MAP-Elites to evolve a diverse set of neural network controllers and find different strategies when playing with one deck. Steckel and Schrum, (2021) use MAP-Elites to illuminate the space for generated levels by different GANs in the Lone Runner game. Wan et al., (2024) apply QD principles to adversarial imitation learning to train agents on several behaviors simultaneously and show that it can potentially improve any inverse RL method. Samvelyan et al., 2024a apply MAP-Elites to explore adversarial scenarios for a pre-trained DRL agent in a multi-agent game environment (Google Research Football), uncovering strategic weaknesses in the agent’s behavior. Samvelyan et al., 2024b use MT-ME to find a set of diverse adversarial safety attacks on an LLM, using an LLM for the variation operator and for evaluating the fitness of the attack. These methods only consider optimizing one side of the adversarial problem, significantly limiting the illumination.
Few QD methods currently try to coevolve both sides. Costa et al., (2020) use NSLC to coevolve GANs (i.e., the generators and descriptors), showing that it improves diversity and discovers better models. Closer to our work, Dharna et al., (2024) propose a self-play QD algorithm to coevolve adversarial controllers as Python code for an asymmetric adversarial game between pursuer and evader agents. They also use the embedding space of a foundation model for diversity, but use an LLM’s embedding space from the Python controller script, which relates to genotypic diversity, not behavioral, as in this paper. Their work also uses an unstructured archive to handle the high-dimensional embedding space. However, they employ an NSLC-type unstructured archive to define novelty, which requires a more problem-specific tuning of the novelty threshold parameter compared to the growing archives used in GAME.
@@
Figure 2: GAME is an adversarial coevolution QD algorithm that iterates the illumination of one side of an adversarial problem using MTMB-ME (Anne and Mouret,, 2023), switching sides at each generation to promote an arms race dynamics that expands the illumination. One key feature of GAME is the ability to use a VEM as a behavior space.
One limitation of MAP-Elites is the need to define a behavior space to determine what is different. Automated Search for Artificial Life (ASAL) (Kumar et al.,, 2024) have shown that a VEM, such as CLIP (Radford et al.,, 2021), can be used to explore the space of artificial life substrates. Inspired by them, we propose, to our knowledge, the first use of a VEM as a behavior space for a MAP-Elites algorithm.
One difficulty with such a space is its high dimensionality (e.g., 2 56025602\,5602 560 dimensions in this paper). The original MAP-Elites uses a grid to divide the behavior space, which is not scalable to high dimensions (as the number of cells grows exponentially with the dimensions). One solution is using Centroidal Voronoi Tessellations (CVT) (Vassiliades et al.,, 2016) to divide the behavior space uniformly. However, in very high-dimensional spaces, one cannot easily predict which proportion of the space will be covered by the studied system, making it difficult to tune the selection pressure. If too many cells are available to fill, the pressure will be too low, reducing quality, and if there are not enough cells, the pressure will be too high, reducing diversity. One solution is to learn a low-dimensional representation using an auto-encoder (Cully,, 2019), which requires extra computation.
Our intuition is that we can grow the archive to cover the relevant part of the behavior space instead of defining it beforehand. This is similar to how NSLC defines novelty with a growing archive of solutions. This is called an unstructured archive (Vassiliades et al.,, 2017). Unlike NSLC, which uses a distance parameter threshold to determine that a new solution should be added to the archive, unstructured archives have a predefined maximal number of cells and only update the position and boundaries of those cells. This can be done by updating the centroids of CVT when the replacement of an old centroid by a new one would increase the minimal distance between all centroids, thus growing the archive.
@@
GAME is a QD algorithm that aims to illuminate both sides (Blue and Red) of an adversarial problem using coevolution to nurture an open-ended discovery of solutions (Fig. 2). It is divided into two levels: (1) Intergenerational illumination of solutions, switching from one side to the other at each generation, Alg. 1 (a, f, and g); and (2) Intragenerational execution of MTMB-ME at each generation, using one side as tasks and the other as solutions, Alg. 2 (b, c, d, and e).
@@
        pick a $task$ at random and use a variation operator on randomly picked elites from the whole archive to generate a candidate solution $s$;
@@
        evaluate $s$ against the $task$, collect the video and fitness $f$, subsample the video (five uniformly spaced frames) to get $frames$, get the embedding ($512$D) from each using a VEM (CLIP), and concatenate to get the behavior descriptor $b$ (2 56025602\,5602 560D);
@@
        update $task$’s archive (Alg. 3): (1) if the new behavior $b$ is farther from all the current cell’s centroids than the closest pair of centroids, it is added to the archive and one of the two centroids from the pair is removed; (2) else if the new fitness $f$ is greater than the corresponding cell’s fitness then the new solution becomes the elite of its cell;
@@
3: $\mathcal{B}=\emptyset$triangleright For storing bootstrapping evaluations
4:for $gen\_id=1$ in $N_{gen}$ do
5:      S $\leftarrow$ $S_{Red}$ if $gen\_id$ is odd else $S_{Blue}$
6:      $\mathcal{A}\leftarrow\operatorname{MTMB-ME}(Tasks,S,\mathcal{B})$triangleright (b-e) - Alg. 2
7:      $Behaviors$leftarrow Aggregate the archive’s elites $\triangleright$ (f)
@@
10:      $Generations[gen\_id]=Tasks$
11:      $\mathcal{B}\leftarrow$ $Tasks$ versus $Generations[gen\_id-1]$triangleright (g)
12: return $Generations$
@@
3:      $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$triangleright Alg. 3
@@
5:      $task\leftarrow$ Select a task at random from $Tasks$triangleright (c)
6:     if  $\mathcal{A}$ has less than $N_{init}$ elites then
@@
10:      $f$, $video\leftarrow\operatorname{Evaluate}(s,task)$triangleright (d)
11:      $frames\leftarrow$ Subsample the $video$
12:      $b$leftarrow Apply the VEM on each frame
13:      $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$triangleright (e) - Alg. 3
@@
When using a VEM, GAME follows ASAL by using cosine similarity to compute the distance between two behaviors (which are 2 56025602\,5602 560D):
@@
The reason is that Euclidean distance loses meaning as the number of dimensions increases. This doesn’t impact the growing archive, as it uses the relative comparison of distances between behaviors and not their actual positions.
@@
One point to consider when growing an archive is that when a new behavior takes over a cell, it can “steal” the elites of neighboring cells as it changes the whole CVT, which creates “holes”. A simple solution to fill those “holes” is to reinstate the centroid as the elite of its cell, as it can never be “stolen.” It only requires storing an additional solution per cell. More elaborate heuristics can be used, e.g., storing all the past elites and checking with the current CVT. However, we found that this doesn’t significantly improve performance for most problems while increasing computation and memory space.
@@
1: $(C,E,E_{backup})=A$triangleright Centroids, Elites and Backup Elites
@@
10:      $d=\min\{dist(b,C_{i})\}_{0\leq i<N_{cells}}$
11:      $c_{id}=\operatorname{find\_cell}(C,b)$triangleright Closest centroid’s index
@@
14:          $d_{j}\leftarrow\min\{dist(C_{j},C_{i})\}_{0\leq i\neq b<N_{cells}}$
15:          $d_{k}\leftarrow\min\{dist(C_{k},C_{i})\}_{0\leq i\neq a<N_{cells}}$
@@
23:     else if f $>$ E[$c_{id}$].f then $\triangleright$ Better fitness
@@
Case study: Multi-agent adversarial game
Figure 3: GAME’s variants and ablations comparisons. The solid line represents the median, and the shaded area shows the min and max of $3$ replications over $20$ generations. (a) Diversity-only and GAME-SO show the largest increase in solution size, but (b) GAME-SO shows the largest increase in complexity. (c-d) Quality-only and GAME-MO (no VEM) lead to the worst performances, while the others are similar. (e) Removing bootstrapping leads to a constant discovery of novel solutions.
To evaluate GAME, we applied it to a competitive RTS game research environment, Parabellum (Anne et al.,, 2025). Competitive games are inherently adversarial, making them a direct target for illuminating adversarial problems. In addition, RTS games employ many agents in a top-down view of the map, and characterizing the multi-agent behavior is not trivial. Parabellum also shares similarities with ALife simulation, with many agents moving and interacting in a 2D map, allowing us to study using a VEM as a behavior space.
@@
Parabellum is a multi-agent game where two sides, Blue and Red, try to eliminate the other, i.e., the fitness of each side is the sum of the depleted health of the opposing side’s units. Each unit can only see other units within its sight range ($15$ m with the map being $100$ m wide). At each time step, each unit can either move, stand, attack an enemy in reach, or heal an ally in reach, given its local observation (position, type, side, and health of all the units in sight). It is implemented in JAX (Bradbury et al.,, 2018), a Python library for just-in-time compilation and vectorization. It permits fast parallelization of the units’ behaviors using a dedicated GPU. Parabellum contains stochasticity, but we use the same seed for all encounters to make comparisons fair.
We defined five types of units (each with a different color for visualization): spearman: slow and high-health close combat unit; archer: low-health long-range combat unit; cavalry: fast and medium health close combat unit; healer: low-health healing unit with close range; grenadier: low-health mid-range unit that inflicts area damage to all units.
Each side comprises $32$ units of each type uniformly spread in each side’s starting area, which are placed so that no unit can initially see an enemy unit (see Fig. 1 frames 1).
@@
More specifically, the units from each side share the same BT, meaning that the search space of GAME is one BT for the Red side and one BT for the Blue side. At each evaluation, each unit visits its BT starting from the root node, and undergoes a left-most depth-first visit of the tree until it finds a valid action. The BT comprises intermediate nodes: Sequence/Failwith nodes, which stop at the first invalid/valid node, and leaves nodes: Action nodes that return an action and its validity, and Condition nodes that return a boolean. To get a fast evaluation, we implemented a BT evaluation function in JAX that allows vectorization of the BT evaluation at the price of setting a maximal number of leaves (100).
Their tree nature allows the use of genetic variation operators for trees. For example, Iovino et al., (2021) evolve BTs for robot control, and Montague et al., (2023) use MAP-Elites with BTs to learn controllers for heterogeneous robot swarms. Taking inspiration from them, GAME randomly selects a variation operator: deleting a sub-tree ($35$%), adding a random node at a random location ($21$%), mutating a node by changing its parameters ($7$%), replacing a node with a random node (7%), or doing a crossover, i.e., copying a random sub-tree of another BT at a random location, ($30$%). Those probabilities could be tuned, but do not seem to change GAME’s performances significantly.
Another advantage of using BT is that it is straightforward to increase its complexity by allowing it to grow, promoting an open-ended evolution (Harrington and Pollack,, 2019). Such an increase in complexity is not trivial to obtain with neural networks and must rely on specific methods such as NEAT (Stanley and Miikkulainen,, 2002).
The BT evolution requires the design of atomic functions (the Actions and Conditions) that take the unit’s local observation and return the corresponding output. To allow synchronization between units, each side can set a target position on the map and move toward it even if it’s out of sight. Different qualifiers parametrize the atomics to improve interpretation and usage, e.g., a target qualifier corresponding to either closest, farthest, weakest, strongest, or random.
The available Actions are: Stand (do nothing); Attack (attack the target enemy in reach of a given type); Heal (heal the target ally in reach of a given type); Move (move toward/away from the target ally/enemy of a given type); Go To (got to the target position); Set Target (mark the position of the target ally/enemy as target position).
The available Conditions are: In Sight (Is there an ally/enemy of a given type in sight?) In Reach (Is there an ally/enemy of a given type in reach?) Is Dying (Is my health or an ally/enemy’s health below a given threshold?) Is Type (Am I of a given type?) Is Set Target (Is the target position set on the map?)
@@
  • GAME-MO: full variant with a multi-objective fitness that first maximizes the opposing side’s depleted health and only secondly minimizes the size of the BT;
  • GAME-MO (no bootstrap): GAME-MO with an ablation of the bootstrapping between generations;

### v3

@@ Abstract
Quality-Diversity (QD) algorithms seek to discover diverse, high-performing solutions across a behavior space, contrasting with conventional optimization methods that target a single optimum. Adversarial problems present unique challenges for QD approaches, as the competing nature of opposing sides creates interdependencies that complicate the evolution process. Existing QD methods applied to such scenarios typically fix one side, constraining behavioral diversity. We present Generational Adversarial MAP-Elites (GAME), a coevolutionary QD algorithm that evolves both sides by alternating which side is evolved at each generation. By integrating a vision embedding model, our approach eliminates the need for domain-specific behavior descriptors and instead operates on video. We validate GAME across three distinct adversarial domains: a multi-agent battle game, a soft-robot wrestling environment, and a deck building game. Our experiments reveal several evolutionary phenomena, including arms-race-like dynamics, enhanced novelty through generational extinction, and the preservation of neutral mutations as crucial stepping stones toward the highest performance. While GAME successfully illuminates all adversarial problems, its capacity for truly open-ended discovery remains constrained by the finite nature of the underlying search spaces. These findings establish GAME’s broad applicability while highlighting opportunities for future research into open-ended adversarial coevolution.
@@ Introduction
Figure 1: GAME’s illumination of a multi-agent adversarial game. The point cloud is a 2D PCA projection (2222% and 1212% explained variance) of the intergenerational tournament between elites found for one run of GAME across 2020 generations with 100 000100\,000 evaluations per generation. We display timed snapshots of eight duels exhibiting different behaviors. We also indicate the fitness of both sides, represented as the percentage of the opposing side’s depleted health. Videos of these duels and supplementary data are available in the repository https://github.com/Timothee-ANNE/GAME.
Quality Diversity (QD) (Pugh et al.,, 2016) is a branch of evolutionary algorithms that seeks to illuminate a problem, i.e., to discover a diverse set of high-quality solutions that maximize a fitness function while covering a behavior space. It has demonstrated success across various domains, including robotics (Cully et al.,, 2015), procedural content generation (Gravina et al.,, 2019), chemical synthesis (Jiang et al.,, 2022), and aeronautics (Brevault and Balesdent,, 2024).
@@
Several such efforts have already been undertaken. For instance, in the turn-based strategy card game Hearthstone (Blizzard Entertainment,, 2014), Fontaine et al., (2019) illuminate the deck space, while Fontaine et al., (2020) evolve a set of neural network controllers with a fixed deck to find different strategies. More recently, Samvelyan et al., 2024a investigate strategic weaknesses in a Deep Reinforcement Learning (DRL) agent during a simulated football game, while Samvelyan et al., 2024b search for different safety-violating prompts for a Large Language Model (LLM). A limitation of these approaches is that they only optimize one side of the adversarial problem. This is problematic, as adversarial systems often lead to an arms race (Dawkins and Krebs,, 1979), which requires improvements on both sides to emerge.
@@
This paper significantly extends our conference paper (Anne et al., 2025a, ), providing a comprehensive introduction to Generational Adversarial MAP-Elites (GAME), a new QD algorithm for adversarial problems, applied to three different adversarial domains: (1) a multi-agent battle game called Parabellum (Anne et al., 2025b, ); (2) a 2D soft-robot custom EvoGym called Wrestling (Bhatia et al.,, 2021); and (3) a deck building game called Hearthbreaker (Daniel et al.,, 2014), which is a Python simulator of Hearthstone (Blizzard Entertainment,, 2014).
Combined with a Vision Embedding Model (VEM) such as CLIP (Radford et al.,, 2021), GAME can operate directly on videos of behaviors, eliminating the need for handcrafted behavior descriptors. GAME handles the high dimensionality of the embedding space by using growing unstructured archives (Vassiliades et al.,, 2017).
We demonstrate through ablation studies in Parabellum that all components of GAME are necessary to illuminate the space of possible strategies effectively. In Wrestling, we show that GAME shows promise for the evolution of artificial creatures, and in Hearthbreaker, we show that GAME’s adversarial coevolution leads to solutions of higher quality but smaller coverage than applying MAP-Elites against a fixed opponent deck.
The paper’s main contributions are:
  • a new QD algorithm, GAME, for adversarial problems illumination;
  • the use of GAME with a VEM as a domain-agnostic behavior space in two distinct domains, a multi-agent battle game and a 2D soft-robot wrestling environment, showing GAME’s generality;
  • the comparison of GAME (without the VEM) against a one-sided illumination in a deck building domain.
Related Work
@@
GAME aims to create an open-ended adversarial coevolution of solutions that broadens illumination. Creating an open-ended artificial environment that continually presents new, interesting challenges and, in turn, fosters the evolution of novel and meaningful solutions to those challenges is one of the main quests in the artificial life community (Bedau et al.,, 2000; Dorin and Stepney,, 2024).
@@
A form of not strictly adversarial problem is the coevolution of an environment and the agent interacting with it. The objective is for the environment to evolve in a way that presents new challenges that are neither too easy nor too difficult, effectively creating a curriculum for the agent by offering a sequence of meaningful stepping stones. Brant and Stanley, (2017) coevolve mazes and maze-solving agents using minimal criterion coevolution, i.e., any agent that solves a maze and any maze that is solved at least once can reproduce. Their approach demonstrates that no handcrafted fitness function or behavior descriptor is required.
@@
Those works are mainly designed to propose an automatic curriculum learning to discover robust agents that can solve challenging environments without explicitly trying to illuminate the space of agents or environments. In contrast, GAME explicitly aims to illuminate both sides of the adversarial problem.
@@
Self-play is a closely related field, which creates a learning curriculum by making the agent play against itself or an earlier version of itself. AlphaStar (Vinyals et al.,, 2019) achieves grandmaster level in the real-time strategy (RTS) player-versus-player game StarCraft 2. It uses a league of agents that serve as training opponents for the main agents. Using a league prevents strategy collapse, improves robustness, and helps avoid local minima and forgetting.
@@
QD algorithms have been used to illuminate adversarial problems. For example, Fontaine et al., (2019) propose MAP-Elites with sliding boundaries to build different decks in Hearthstone (Blizzard Entertainment,, 2014), using as opponents first a fixed set of decks and then the best decks found in the first phase to explore the possible counters. Using the six best found decks as opponents, Fontaine et al., (2020) use Covariance Matrix Adaptation MAP-Elites to evolve a diverse set of neural network controllers and find different strategies when playing with one deck. Steckel and Schrum, (2021) use MAP-Elites to illuminate the space for generated levels by different GANs in the Lone Runner game. Wan et al., (2024) apply QD principles to adversarial imitation learning to train agents on several behaviors simultaneously and show that it can potentially improve any inverse RL method. Samvelyan et al., 2024a apply MAP-Elites to explore adversarial scenarios for a pre-trained DRL agent in a multi-agent game environment (Google Research Football), uncovering strategic weaknesses in the agent’s behavior. Samvelyan et al., 2024b use MT-ME to find a set of diverse adversarial safety attacks on an LLM, using an LLM for the variation operator and for evaluating the fitness of the attack. These methods only consider optimizing one side of the adversarial problem, significantly limiting the illumination. The third case study, Hearthbreaker, is an initial attempt to compare GAME’s ability to alternate the illumination against a one-sided illumination similar to Fontaine et al., (2019). A noteworthy difference is that we use the Hearthbreaker (Daniel et al.,, 2014) simulator in Python, while Fontaine et al., (2019) use the SabberStone (Milva et al.,, 2016) simulator in C#, which proposes better AI agents for playing the deck.
Few QD methods currently try to coevolve both sides. Costa et al., (2020) use NSLC to coevolve GANs (i.e., the generators and descriptors), showing that it improves diversity and discovers better models. Closer to our work, Dharna et al., (2024) propose a self-play QD algorithm to coevolve adversarial controllers as Python code for an asymmetric adversarial game between pursuer and evader agents. They also use the embedding space of a foundation model for diversity, but use an LLM embedding space of the Python controller script, which relates to genotypic diversity rather than behavioral diversity, as in this paper. Their work also uses an unstructured archive to handle the high-dimensional embedding space. However, they employ an NSLC-type unstructured archive to define novelty, which requires more problem-specific tuning of the novelty threshold parameter compared to the growing archives used in GAME.
@@
Figure 2: GAME is an adversarial coevolution QD algorithm that iterates the illumination of one side of an adversarial problem using MTMB-ME (Anne and Mouret,, 2023), switching sides at each generation to promote arms-race dynamics that expand the illumination. One key feature of GAME is the ability to use a VEM as a behavior space.
One limitation of MAP-Elites is the need to define a behavior space to determine what is different. Automated Search for Artificial Life (ASAL) (Kumar et al.,, 2024) has shown that a VEM, such as CLIP (Radford et al.,, 2021), can be used to explore the space of artificial life substrates. Inspired by them, we propose, to our knowledge, the first use of a VEM as a behavior space for a MAP-Elites algorithm.
One difficulty with such a space is its high dimensionality (e.g., 2 5602\,560 dimensions in this paper). The original MAP-Elites uses a grid to divide the behavior space, which is not scalable to high dimensions (as the number of cells grows exponentially with the dimensions). One solution is using Centroidal Voronoi Tessellations (CVT) (Vassiliades et al.,, 2016) to divide the behavior space uniformly. However, in very high-dimensional spaces, one cannot easily predict what proportion of the space will be covered by the studied system, making it difficult to tune the selection pressure. If too many cells are available to fill, the pressure will be too low, reducing quality, and if there are not enough cells, the pressure will be too high, reducing diversity. One solution is to learn a low-dimensional representation using an autoencoder (Cully,, 2019), which requires extra computation.
Our intuition is that we can grow the archive to cover the relevant part of the behavior space instead of defining it beforehand. This is similar to how NSLC defines novelty with a growing archive of solutions. This is called an unstructured archive (Vassiliades et al.,, 2017). Unlike NSLC, which uses a distance-threshold parameter to determine whether a new solution should be added to the archive, unstructured archives have a predefined maximal number of cells and only update the position and boundaries of those cells. This can be done by updating the centroids of the CVT when the replacement of an old centroid by a new one would increase the minimal distance between all centroids, thus growing the archive.
@@
GAME is a QD algorithm that aims to illuminate both sides (Blue and Red) of an adversarial problem using coevolution to nurture open-ended discovery of solutions (Fig. 2). It is divided into two levels: (1) intergenerational illumination of solutions, switching from one side to the other at each generation, Alg. 1 (a, f, and g); and (2) intragenerational execution of MTMB-ME at each generation, using one side as tasks and the other as solutions, Alg. 2 (b, c, d, and e).
@@
        pick a tasktask at random and use a variation operator on randomly picked elites from the whole archive to generate a candidate solution $s$;
@@
        evaluate $s$ against the tasktask, collect the video and fitness $f$, subsample the video (five uniformly spaced frames) to get framesframes, obtain the embedding (512512D) from each using a VEM (CLIP), and concatenate to get the behavior descriptor $b$ (2 5602\,560D);
@@
        update tasktask’s archive (Alg. 3): (1) if the new behavior $b$ is farther from all the current cells’ centroids than the closest pair of centroids, it is added to the archive and one of the two centroids from the pair is removed; (2) else if the new fitness $f$ is greater than the corresponding cell’s fitness, then the new solution becomes the elite of its cell;
@@
3: $\mathcal{B}=\emptyset$ $\triangleright$ For storing bootstrapping evaluations
4:for gen_id=1gen\_id=1 in $N_{gen}$ do
5:   S $\leftarrow$ $S_{Red}$ if gen_idgen\_id is odd else $S_{Blue}$
6:   $\mathcal{A}\leftarrow\operatorname{MTMB-ME}(Tasks,S,\mathcal{B})$ $\triangleright$ (b-e) - Alg. 2
7:   BehaviorsBehaviors $\leftarrow$ Aggregate the archive’s elites $\triangleright$ (f)
@@
10:   Generations[gen_id]=TasksGenerations[gen\_id]=Tasks
11:   $\mathcal{B}\leftarrow$ TasksTasks versus Generations[gen_id1]Generations[gen\_id-1] $\triangleright$ (g)
12: return GenerationsGenerations
@@
3:   $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$ $\triangleright$ Alg. 3
@@
5:   $task\leftarrow$ Select a task at random from TasksTasks $\triangleright$ (c)
6:  if $\mathcal{A}$ has fewer than $N_{init}$ elites then
@@
10:   $f$, $video\leftarrow\operatorname{Evaluate}(s,task)$ $\triangleright$ (d)
11:   $frames\leftarrow$ Subsample the videovideo
12:   $b$ $\leftarrow$ Apply the VEM on each frame
13:   $\mathcal{A}\leftarrow\operatorname{Update}(\mathcal{A}[task],s,f,b)$ $\triangleright$ (e) - Alg. 3
@@
When using a VEM, GAME follows ASAL by using cosine similarity to compute the distance between two behaviors (which are 2 5602\,560D):
@@
The reason is that Euclidean distance loses meaning as the number of dimensions increases. This does not impact the growing archive, as it uses relative comparisons of distances between behaviors rather than their actual positions.
@@
One point to consider when growing an archive is that when a new behavior takes over a cell, it can “steal” the elites of neighboring cells as it changes the whole CVT, which creates “holes.” A simple solution to fill those holes is to reinstate the centroid as the elite of its cell, as it can never be “stolen.” It only requires storing an additional solution per cell. More elaborate heuristics can be used, e.g., storing all the past elites and checking with the current CVT. However, we found that this does not significantly improve performance for most problems while increasing computation and memory usage.
@@
1: $(C,E,E_{backup})=A$ $\triangleright$ Centroids, Elites, and Backup Elites
@@
10:   $d=\min\{\operatorname{dist}(b,C_{i})\}_{0\leq i<N_{cells}}$
11:   $c_{id}=\operatorname{find\_cell}(C,b)$ $\triangleright$ Closest centroid’s index
@@
14:    $d_{j}\leftarrow\min\{\operatorname{dist}(C_{j},C_{i})\}_{0\leq i\neq j<N_{cells}}$
15:    $d_{k}\leftarrow\min\{\operatorname{dist}(C_{k},C_{i})\}_{0\leq i\neq k<N_{cells}}$
@@
23:  else if f >> E[$c_{id}$].f then $\triangleright$ Better fitness
@@
Case study: Multi-agent adversarial game, Parabellum
Figure 3: GAME’s variants and ablations comparisons. The solid line represents the median, and the shaded area shows the min and max of 33 replications over 2020 generations. (a) Quality-only and Diversity-only show the largest increase in solution size, but (b) Quality-only and GAME-SO show the largest increase in complexity. (c–d) Quality-only and GAME-MO (no VEM) lead to the worst QD performances. (e) Removing bootstrapping leads to a constant discovery of novel solutions.
To evaluate GAME, we first applied it to an adversarial battle game research environment, Parabellum (Anne et al., 2025b, ). Battle games are inherently adversarial, making them a direct target for illuminating adversarial problems. In addition, battle games employ many agents in a top-down view of the map, and characterizing multi-agent behavior is not trivial. Parabellum also shares similarities with ALife simulations, where many agents move and interact on a 2D map, which inspired us to apply the same VEM as in ASAL (Kumar et al.,, 2024).
@@
Parabellum is a multi-agent battle game where two sides, Blue and Red, try to eliminate the other, i.e., the fitness of each side is the sum of the depleted health of the opposing side’s units. Each unit can only see other units within its sight range ($15$ m, with the map being $100$ m wide). At each time step, each unit can either move, stand, attack an enemy in reach, or heal an ally in reach, given its local observation (position, type, side, and health of all the units in sight). It is implemented in JAX (Bradbury et al.,, 2018), a Python library for just-in-time compilation and vectorization. It permits fast parallelization of the units’ behaviors using a dedicated GPU. Parabellum contains stochasticity, but we use the same seed for all encounters to make comparisons fair.
We defined five types of units (each with a different color for visualization): spearman: slow and high-health close combat unit; archer: low-health long-range combat unit; cavalry: fast and medium-health close combat unit; healer: low-health healing unit with close range; grenadier: low-health mid-range unit that inflicts area damage to all units.
Each side comprises $32$ units of each type uniformly spread in each side’s starting area, which are placed so that no unit initially sees an enemy unit (see Fig. 1 frames 1).
@@
More specifically, the units from each side share the same BT, meaning that the search space of GAME is one BT for the Red side and one BT for the Blue side. At each evaluation, each unit visits its BT starting from the root node and undergoes a left-most depth-first visit of the tree until it finds a valid action. The BT comprises intermediate nodes: Sequence/Failwith nodes, which stop at the first invalid/valid node, and leaf nodes: Action nodes that return an action and its validity, and Condition nodes that return a boolean. To obtain fast evaluations, we implemented a BT evaluation function in JAX that allows vectorization of the BT evaluation at the price of setting a maximal number of leaves (100).
Their tree nature allows the use of genetic variation operators for trees. For example, Iovino et al., (2021) evolve BTs for robot control, and Montague et al., (2023) use MAP-Elites with BTs to learn controllers for heterogeneous robot swarms. Taking inspiration from them, GAME randomly selects a variation operator: deleting a subtree (3535%), adding a random node at a random location (2121%), mutating a node by changing its parameters (77%), replacing a node with a random node (77%), or performing a crossover, i.e., copying a random subtree of another BT at a random location (3030%). Those probabilities could be tuned but do not seem to change GAME’s performance significantly.
Another advantage of using BTs is that it is straightforward to increase their complexity by allowing them to grow, promoting open-ended evolution (Harrington and Pollack,, 2019). Such an increase in complexity is not trivial to obtain with neural networks and must rely on specific methods such as NEAT (Stanley and Miikkulainen,, 2002).
The BT evolution requires the design of atomic functions (the Actions and Conditions) that take the unit’s local observation and return the corresponding output. To allow synchronization between units, each side can set a target position on the map and move toward it even if it is out of sight. Different qualifiers parametrize the atomics to improve interpretation and usage, e.g., a target qualifier corresponding to either closest, farthest, weakest, strongest, or random.
The available Actions are: Stand (do nothing); Attack (attack the target enemy in reach of a given type); Heal (heal the target ally in reach of a given type); Move (move toward/away from the target ally/enemy of a given type); Go To (go to the target position); Set Target (mark the position of the target ally/enemy as target position).
The available Conditions are: In Sight (Is there an ally/enemy of a given type in sight?) In Reach (Is there an ally/enemy of a given type in reach?) Is Dying (Is my health or an ally’s/enemy’s health below a given threshold?) Is Type (Am I of a given type?) Is Set Target (Is the target position set on the map?)